he slope of the AE curve is .80. What is the multiplier? Everything else the same, by how much 

does equilibrium aggregate expenditure increase if

a) exports increase from $1.75 trillion to $2.25 trillion. (8 marks)

b) government expenditure on goods and services decrease from $2.0 trillion to $1.8 trillion.

(6 marks)

c) investment increases from $1.2 trillion to $2.3 trillion. (6 marks)


1
Expert's answer
2021-09-14T07:29:44-0400

Solution:

The slope of AE curve is the MPC = 0.80

Multiplier = "\frac{1}{1 - MPC} = \frac{1}{1 - 0.8} = \frac{1}{0.2} = 5"


The Multiplier = 5

 

a.). "\triangle" Exports = 2.25 – 1.75 = 0.5

"\triangle AE = \triangle Exports\times Multiplier = 0.5\times 5 = 2.5"

The equilibrium aggregate expenditure will increase by $2.5 trillion dollars.

 

b.). "\triangle" G = 1.8 – 2.0 = -0.2

"\triangle AE = \triangle G\times Multiplier = -0.2\times 5 = -1"

The equilibrium aggregate expenditure will decrease by $1 trillion dollars.

 

c.). "\triangle" I= 2.3 – 1.2 = 1.1

"\triangle AE = \triangle I\times Multiplier = 1.1\times 5 = 5.5"

The equilibrium aggregate expenditure will increase by $5.5 trillion dollars.